OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. Which recursive equation represents the pattern? Please help, I don't understand!!!

OpenStudy (ashleyisakitty):

a4 =a3+5=(4+5)=9 the answer is d.

OpenStudy (anonymous):

an = an – 1 + 3n – 2 an = an – 1 + 2n – 3 an = an – 1 + (3n – 2) an = an – 1 + (2n – 3) These are the answer choices

OpenStudy (ashleyisakitty):

Yes, i know. Look at what i plugged in.

OpenStudy (anonymous):

Thank you!!!